Suppose a car waiting at red traffic light when the light turns to green the car accelerates uniformly at 2m/s within 6 sec then proceeds moving with a uniform speed at the movement the car starts at the green traffic light a truck passes by heading in the direction with a uniform speed of 10m/s. How long and how far from traffic light the car catches up with the truck?


1
Expert's answer
2020-01-23T06:10:09-0500
"x=0.5at^2+(at)(T-t), x=vT"

"0.5at^2+(at)(T-t)=vT"

"0.5(2)(6)^2+(2)(6)(T-6)=10T"

The time will be

"T=18\ s"

The distance:


"d=vT=(10)(18)=180\ m"